# SUMO: Subspace-Aware Moment-Orthogonalization for Accelerating Memory-Efficient LLM Training

### v1

@@ 3.1 Theoretical Motivation: Exact moments orthogonalization leads to significantly faster convergence
Recall that in the low-rank setting, accurately computing the pseudoinverse using singular value decomposition (SVD) is numerically advantageous and reasonably computationally affordable compared to iterative methods such as Newton–Schulz. For a general matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$, where without the loss of generality $m\geq n$, the SVD provides a decomposition $\mathbf{A}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top}$, with $\mathbf{U}\in\mathbb{R}^{m\times m}$, $\mathbf{\Sigma}\in\mathbb{R}^{m\times n}$, and $\mathbf{V}\in\mathbb{R}^{n\times n}$. The Moore–Penrose pseudoinverse is then calculated as $\mathbf{A}^{\dagger}=\mathbf{V}\mathbf{\Sigma}^{\dagger}\mathbf{U}^{\top}$, requiring approximately $4nm^{2}+8n^{3}$ floating-point operations (FLOPs) for the initial decomposition, and an additional $mn^{2}+m^{2}n$ FLOPs for subsequent multiplications, totaling roughly $4mn^{2}+8n^{3}+nm^{2}+n^{2}m$ FLOPs.
Alternatively, approximating the inverse of $\mathbf{A}^{\top}\mathbf{A}\in\mathbb{R}^{m\times m}$ using Newton–Schulz iterations involves $nm^{2}$ FLOPs to form the matrix $\mathbf{A}^{\top}\mathbf{A}$, approximately $20m^{3}+10m^{2}$ FLOPs for five iterations, and an additional $m^{2}n$ FLOPs to multiply by $\mathbf{A}^{\top}$, resulting in a total of about $nm^{2}+m^{2}n+20m^{3}+10m^{2}$ FLOPs. For example, when the rank is $m=8$ and $n=1024$, the SVD approach requires approximately twice the number of operations compared to Newton–Schulz5. Nonetheless, given the superior numerical stability and inherent optimality of the SVD-based method, this moderate increase in computational effort remains acceptable, especially when accuracy and stability are prioritized.
@@ 3.2 Method
We are now ready to present our main algorithm designed to accelerate the low-rank optimization scheme outlined in Algorithm 1. A detailed mathematical formulation of the weight update rule proposed in this paper can be found in Appendix C. The algorithm consists of four primary blocks, all contained within an outer loop that continues until convergence is achieved or a predefined number of epochs is reached. Each block serves a specific purpose, which will be explained in detail below.
@@
        $\mathbf{Q}_{t}\leftarrow\operatorname{Truncated\_Randomized\_SVD}(\mathbf{G}_{t})$ Alternatively $\operatorname{Truncated\_SVD}(\mathbf{G}_{t})$
@@
     if $\frac{\left\|O^{(t)}\right\|}{\left\|O^{(t)}\right\|}>\gamma$ then $\mathbf{O}^{(t)}\leftarrow\frac{\mathbf{O}^{(t)}}{\left\|\mathbf{O}^{(t)}\right\|}\cdot\gamma\left\|\mathbf{O}^{(t-1)}\right\|$
@@
     ${\bf W}^{(t)}\longleftarrow{\bf W}^{(t-1)}-\alpha\eta{{\bf Q}}^{(t)}{\bf O}^{(t)}-\eta\cdot\lambda{\bf W}^{(t-1)}$
@@
  • Block 1: We select the subspace along the directions of the $r$ largest eigenvectors, but since computing full SVD for large matrices is computationally intensive and memory-demanding, we leverage the Randomized-SVD by [44], which is an efficient technique for producing a “good" proxy for the optimal low-rank approximation. It solves the optimization problem $\arg\min_{\mathbf{Q}\in\mathbb{R}^{n\times r}}\left\|{\bf G}-\mathbf{QQ}^{\top}{\bf G}\right\|_{F},$ and approximates the matrix ${\bf G}$ as ${\bf G}_{\mathsf{app},r}\approx\mathbf{QQ}^{\top}{\bf G},$ that requires $O(mnr+mr^{2})$ operations, instead of $O(\min(mn^{2},m^{2}n))$ applied by SVD.
  • Block 1.1: We transform the first-order moments evaluated during the low-rank optimization steps, which occur in Block 2, between the preceding and the newly updated subspace. This transformation is required because, as will be demonstrated later, within Block 2, the first moments of the gradients are aligned with the previously projected subspace. Consequently, a transformation is necessary to translate them from the former subspace to the current one.
  • Block 2: Here we calculate the (steepest) optimization step. SVD operation is adopted to solve exactly $({\bf M}^{(t)}{{\bf M}^{(t)}}^{\top})^{-1/2}{\bf M}^{(t)}$. Let ${\bf U}\boldsymbol{\Sigma}\mathbf{V}^{\top}=\hat{{\bf M}}^{(t)}$ be the singular value decomposition (SVD) of $\hat{{\bf M}}^{(t)}$, we will have $({\bf M}^{(t)}{{\bf M}^{(t)}}^{\top})^{-1/2}{\bf M}^{(t)}=\mathbf{U}\mathbf{V}^{\mathbf{T}}$, which orthogonalizes $\hat{{\bf M}}^{(t)}$. Formally, the
[…]
  • Block 3: Rather than using standard gradient clipping, we adopt the Norm-growth Limiter (NL) introduced by [21], which has been shown to slightly outperform traditional clipping techniques by better constraining the progression of gradient magnitudes. Specifically, the gradient update is modified as follows, if ${\small{\text{if }\frac{\left\|{\bf O}^{(t)}\right\|}{\left\|{\bf O}^{(t-1)}\right\|}>\gamma\text{ then }{\bf O}^{(t)}\leftarrow\frac{{\bf O}^{(t)}}{\left\|{\bf O}^{(t)}\right\|}\cdot\gamma\left\|{\bf O}^{(t-1)}\right\|,}}$ where the scalar $\gamma$ serves as a growth threshold to regulate abrupt increases in gradient norm from one iteration to the next. We use $\gamma=1.1$, which empirically leads to the highest results.
  • Block 4: The pre-trained model parameters are updated using the low-dimensional back-projection matrix along with weight decay. To ensure stable training across parameter matrices of different shapes, we interpret the root mean square (RMS) magnitude of updates as implicit layer-wise learning rate adaptation, following the approach in [39]. By scaling updates with $\sqrt{\max(m,n)}$, our method compensates for shape-induced differences in magnitude, achieving consistent effective learning rates across layers, similar to adaptive optimizers like AdamW.
For clarity, we can assume, without loss of generality, that $m\geq n$. In the opposite scenario, the projection matrix would multiply the gradient from the right-hand side.
@@
To reduce memory consumption, Algorithm 1 applies per-layer weight updates during backpropagation, following recent works such as [45]. This contrasts with conventional optimizers that store full gradients and update all weights afterward, leading to potential inefficiencies. Details for post-hoc adapter extraction are discussed in Appendix B.
@@ 4 Experiments
[…]
@@
Our model was evaluated using the GLUE benchmark [43] through the fine-tuning of the pre-trained Roberta-base model [42] across eight tasks. The comparative analysis includes full fine-tuning, LoRA, and GaLore methodologies, with the results enumerated in Table 2. The reported metrics are the overall accuracy (matched and mismatched) for MNLI, Matthew’s correlation for CoLA, Pearson correlation for STS-B, F1-score for MRPC, and accuracy for the remaining tasks. Evidently, our approach enhances the fine-tuning accuracy while requiring less training memory, utilizing only a single moment compared to GaLore. The experiments were carried out using the NVIDIA A100 GPU.
@@
To demonstrate the benefits of our method for pre-training LLAMA by following the evaluation from [1]. Specifically, we compare the performance of SUMO to state-of-the-art methods in terms of perplexity and memory efficiency. For this evaluation, we trained large LLaMA-based models on the C4 dataset, a curated and extensive version of the Common Crawl web corpus [46]. This dataset is widely used for pre-training language models and developing word representations. To better reflect real-world pre-training scenarios, we conducted training on a non-repeating, large-scale dataset and scaled model sizes up to 350 million parameters. The results of these experiments are shown in Table 3. Experiments were conducted using an NVIDIA H200 GPU.
Table 3: Comparison of state-of-the-art low-rank algorithms for pre-training LLaMA models of varying sizes on the C4 dataset. The results are reported in terms of validation perplexity. As can be seen, SUMO leads to improved performance with a substantial memory reduction compared to leading parameter-efficient fine-tuning schemes.
@@
Full-Rank $34.06$ ($0.36$G) $25.08$ ($0.76$G) $18.80$ ($2.06$G) $15.56(7.80\mathrm{G})$
GaLore $34.88$ ($0.24$G) $25.36$ ($0.52$G) $18.95$ ($1.22$G) $15.64(4.38\mathrm{G})$
Low-Rank $78.18$ ($0.26$G) $45.51$ ($0.54$G) $37.41$ ($\mathbf{1.08}$G) $142.53(3.57\mathrm{G})$
LoRA $34.99$ ($0.36$G) $33.92$ ($0.80$G) $25.58$ ($1.76$G) $19.21(6.17\mathrm{G})$
ReLoRA $37.04$ ($0.36$G) $29.37$ ($0.80$G) $29.08$ ($1.76$G) $18.33(6.17\mathrm{G})$
SUMO 34.26 ($\mathbf{0.23}$G) 24.87 ($\mathbf{0.51}$G) 18.69 ($1.16$G) 14.68 ($1.16$G)
Training Tokens $1.1$B $2.2$B $6.4$B 13.1B
@@
To evaluate our method’s performance on a complex reasoning task, we use the GSM8K dataset [47], testing systematic generalization. For these experiments, we used a batch size of 32 and 10 epochs for fine-tuning. We present the performance result in Table 5 training Phi-2 (2.7B) model [48], and in Table 5 training LLaMA (3B) model [49]. The results demonstrate that the proposed method significantly improves generalization to out-of-distribution data. The experiments were conducted on an NVIDIA H200 GPU.
@@ 5 Discussion
Empirically, SUMO outperforms prior low-rank methods in both fine-tuning and pretraining tasks, achieving better performance in memory consumption reduction compared to the memory-efficient benchmarks such as Galore. Our theoretical analysis further confirms its superior convergence properties under practical conditions. These findings position SUMO as a simple yet effective alternative to approximate geometric optimizers. Future work may investigate parallel computations for orthogonalization, integrate quantization techniques, and assess the effectiveness of the method in knowledge editing [50] or domain generalization [51, 52].

### v2

@@ 3.1 Theoretical Motivation: Exact moments orthogonalization leads to significantly faster convergence
Recall that in the low-rank setting, accurately computing the pseudoinverse using singular value decomposition (SVD) is numerically advantageous and reasonably computationally affordable compared to iterative methods such as Newton–Schulz. For a general matrix $\mathbf{A}\in\mathbb{R}^{n\times m}$, the SVD provides a decomposition $\mathbf{A}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top}$, with $\mathbf{U}\in\mathbb{R}^{n\times n}$, $\mathbf{\Sigma}\in\mathbb{R}^{n\times m}$, and $\mathbf{V}\in\mathbb{R}^{m\times m}$. The Moore–Penrose pseudoinverse is then calculated as $\mathbf{A}^{\dagger}=\mathbf{V}\mathbf{\Sigma}^{\dagger}\mathbf{U}^{\top}$, requiring approximately $4nm^{2}+8m^{3}$ floating-point operations (FLOPs) for the initial decomposition, and an additional $mn^{2}+m^{2}n$ FLOPs for subsequent multiplications, totaling roughly $4nm^{2}+8m^{3}+mn^{2}+m^{2}n$ FLOPs.
Alternatively, approximating the inverse of $\mathbf{A}^{\top}\mathbf{A}\in\mathbb{R}^{m\times m}$ using Newton–Schulz iterations involves $nm^{2}$ FLOPs to form the matrix $\mathbf{A}^{\top}\mathbf{A}$, approximately $20m^{3}+10m^{2}$ FLOPs for five iterations, and an additional $m^{2}n$ FLOPs to multiply by $\mathbf{A}^{\top}$, resulting in a total of about $nm^{2}+m^{2}n+20m^{3}+10m^{2}$ FLOPs. For example, when the rank is $m=8$ and $n=1024$, the SVD approach requires approximately twice as many operations as Newton–Schulz5. Nonetheless, given the superior numerical stability and inherent optimality of the SVD-based method, this moderate increase in computational effort remains acceptable, especially when accuracy and stability are prioritized.
@@ 3.2 Method
We are now ready to present our main algorithm designed to accelerate the low-rank optimization scheme outlined in Algorithm 1. A detailed mathematical formulation of the weight update rule proposed in this paper is given in Appendix C. The algorithm consists of four primary blocks, all contained within an outer loop that continues until convergence is achieved or a predefined number of epochs is reached. Each block serves a specific purpose, which will be explained in detail below.
  • Block 1 (Adaptive Subspace Selection): We select the subspace along the directions of the $r$ largest eigenvectors, but since computing full SVD for large matrices is computationally intensive and memory-demanding, we leverage the Randomized-SVD by [44], which is an efficient technique for producing a “good" proxy for the optimal low-rank approximation. It solves the optimization problem $\arg\min_{\mathbf{Q}^{(t)}\in\mathbb{R}^{m\times r}}\left\|{\bf G}-\mathbf{Q}^{(t)}{\mathbf{Q}^{(t)}}^{\top}{\bf G}{(t)}\right\|_{F},$ and approximates the matrix $\bf G$} as $\hat{{\bf G}}\approx\mathbf{Q}^{(t)}{\mathbf{Q}^{(t)}}^{\top}{\bf G}{(t)},$ that requires $O(mnr+mr^{2})$ operations, instead of $O(\min(mn^{2},m^{2}n))$ applied by SVD.
  • Block 1.1 (Moment Subspace Transformation): We transform the first-order moments evaluated during the low-rank optimization steps, which occur in Block 2, between the preceding and the newly updated subspace. This transformation is required because, as will be demonstrated later, the first moments of the gradients in Block 2 are aligned with the previously projected subspace. Consequently, a transformation is necessary to translate them from the former subspace to the current one.
  • Block 2 (Low-Rank Steepest Descent Optimization): Here we calculate the (steepest) optimization step. SVD operation is adopted to solve exactly $({\bf M}^{(t)}{{\bf M}^{(t)}}^{\top})^{-1/2}{\bf M}^{(t)}$. Let ${\bf U}\boldsymbol{\Sigma}\mathbf{V}^{\top}=\hat{{\bf M}}^{(t)}$ be the singular value decomposition (SVD) of $\hat{{\bf M}}^{(t)}$, we will have $({\bf M}^{(t)}{{\bf M}^{(t)}}^{\top})^{-1/2}{\bf M}^{(t)}=\mathbf{U}\mathbf{V}^{\mathbf{T}}$, which orthogonalizes $\hat{{\bf M}}^{(t)}$. Formally, the
    $\text{Orthogonalization\_SVD}({\bf A})=\underset{{\bf O}}{\operatorname{arg\,min}}\{\|{\bf O}-{\bf A}\|_{F}:\text{either }{\bf O}^{T}{\bf O}={\bf I}\text{ or }{\bf O}{\bf O}^{T}={\bf I}\}.$
  • Block 3: Rather than using standard gradient clipping, we adopt the Norm-growth Limiter (NL) introduced in [21], which has been shown to slightly outperform traditional clipping techniques by better constraining the progression of gradient magnitudes. Specifically, the gradient update is modified as follows, $\small{\text{if }\frac{\left\|{\bf O}^{(t)}\right\|}{\left\|{\bf O}^{(t-1)}\right\|}>\gamma\text{ then }{\bf O}^{(t)}\leftarrow\frac{{\bf O}^{(t)}}{\left\|{\bf O}^{(t)}\right\|}\cdot\gamma\left\|{\bf O}^{(t-1)}\right\|,}$} where the scalar $\gamma$ serves as a growth threshold to regulate abrupt increases in gradient norm from one iteration to the next. We use $\gamma=1.1$, which empirically yields the best results.
  • Block 4 (Update Step in the Original Space). To better exploit already computed gradient information, we suggest to use the orthogonal term of the gradient that lies outside the low-rank subspace spanned ${\bf O}^{(t)},$ namely ${{\bf G}^{(t)}}^{\perp}={\bf G}^{(t)}-{{{\bf Q}}^{(t)}}\underbrace{{{{\bf Q}}^{(t)}}^{\top}{\bf G}^{(t)}}_{\hat{{\bf G}}^{(t)}}$. Importantly, ${\bf G}^{(t)}$$^{\perp}$ does not interfere with the orthogonalized moment estimation ${\bf O}^{(t)}$, meaning it can be incorporated without compromising stability. Since ${\bf G}^{(t)}$ is already computed and stored in each iteration, no extra memory is required. Furthermore, because ${{\bf Q}}^{(t)}$ is of a low rank (typically rank $4$, 88, or $16$), the additional computationations overhead is negligible. For efficient memory usage, instead of explicitly forming the perpendicular part ${\bf G}^{(t)\perp}$, we work only in the rank-$r$ subspace spanned by ${{\bf Q}}^{(t)}$}. Practically, we use fact that
    $\underbrace{\bigl({\bf G}^{(t)}-{{\bf Q}}^{(t)}{{{\bf Q}}^{(t)}}^{\top}{\bf G}^{(t)}\bigr)}_{{\bf G}^{(t)\perp}}\;+\;{{\bf Q}}^{(t)}{\bf O}^{(t)}={\bf G}^{(t)}-{{\bf Q}}^{(t)}\!\bigl(\hat{{\bf G}}^{(t)}-{\bf O}^{(t)}\bigr),$
    which utilizes the already calculated projected gradient $\hat{{\bf G}}^{(t)}:=\;{{{\bf Q}}^{(t)}}^{\top}{\bf G}^{(t)}$. Accordingly, the pre-trained model parameters are updated along with weight decay,
    ${\bf W}^{(t)}\;\leftarrow\;{\bf W}^{(t-1)}\;-\;\alpha\eta\,\Bigl({\bf G}^{(t)}-{{\bf Q}}^{(t)}\bigl(\hat{{\bf G}}^{(t)}-{\bf O}^{(t)}\bigr)\Bigr)\;-\;\eta\,\lambda\,{\bf W}^{(t-1)}.$
    To ensure stable training across parameter matrices of different shapes, we interpret the root mean square (RMS) magnitude of updates as implicit layer-wise learning rate adaptation, following the approach in [39]. By scaling updates by $\sqrt{\max(m,n)}$, our method compensates for shape-induced magnitude differences, achieving consistent effective learning rates across layers, similar to adaptive optimizers like AdamW.
@@
   $\mathbf{Q}^{(t)}\leftarrow\operatorname{Truncated\_Randomized\_SVD}(\mathbf{G}^{(t)})$ # Alternatively $\operatorname{Truncated\_SVD}(\mathbf{G}^{(t)})$
@@
  if $\frac{\left\|{\bf O}^{(t)}\right\|}{\left\|{\bf O}^{(t)}\right\|}>\gamma$ then $\mathbf{O}^{(t)}\leftarrow\frac{\mathbf{O}^{(t)}}{\left\|\mathbf{O}^{(t)}\right\|}\cdot\gamma\left\|\mathbf{O}^{(t-1)}\right\|$
@@
  ${\bf W}^{(t)}\;\leftarrow\;{\bf W}^{(t-1)}\;-\;\alpha\eta\,\Bigl({\bf G}^{(t)}-{{\bf Q}}^{(t)}\bigl(\hat{{\bf G}}^{(t)}-{\bf O}^{(t)}\bigr)\Bigr)\;-\;\eta\,\lambda\,{\bf W}^{(t-1)}$
@@
Note that, for clarity, we can assume, without loss of generality, that $m\geq n$. In the opposite scenario, the projection matrix would multiply the gradient from the right-hand side.
@@
To reduce memory consumption, Algorithm 1 applies per-layer weight updates during backpropagation, following recent works such as [45]. This contrasts with conventional optimizers, which store full gradients and update all weights afterward, potentially leading to inefficiencies. Details for post-hoc adapter extraction are discussed in Appendix B.
@@ 4 Experiments
Figure 2: SUMO with SVD demonstrates superior convergence speed ($\sim\!1.6\times$ faster), attaining comparable or higher accuracy than GaLore and SUMO with Newton-Schultz5 with significantly fewer optimization steps on QNLI.
@@
Our model was evaluated using the GLUE benchmark [43] through the fine-tuning of the pre-trained Roberta-base model [42] across eight tasks. The comparative analysis includes full fine-tuning, LoRA, and GaLore methodologies, with the results enumerated in Table 2. The reported metrics are the overall accuracy (matched and mismatched) for MNLI, Matthew’s correlation for CoLA, Pearson correlation for STS-B, F1-score for MRPC, and accuracy for the remaining tasks. Evidently, our approach improves fine-tuning accuracy while requiring less training memory, using only a single moment compared to GaLore. The experiments were carried out using the NVIDIA A100 GPU.
@@
To highlight the effectiveness of our method in pre-training, we pre-train LLaMA models following the evaluation protocol of [1], comparing the performance to the state-of-the-art method, in terms of perplexity and memory usage. Specifically, we compare SUMO’s performance with state-of-the-art methods on perplexity and memory efficiency. For this evaluation, we trained large LLaMA-based models on the C4 dataset, a curated and extensive version of the Common Crawl web corpus [46]. This dataset is widely used for pre-training language models and developing word representations. To better reflect real-world pre-training scenarios, we trained on a non-repeating, large-scale dataset and scaled model sizes up to 1 billion parameters. The results of these experiments are shown in Table 3. Experiments were conducted using an NVIDIA H200 GPU.
Table 3: Comparison of state-of-the-art low-rank algorithms for pre-training LLaMA models of varying sizes on the C4 dataset. The results are reported in terms of validation perplexity. As shown, SUMO leads to improved performance with substantial memory reduction compared to leading parameter-efficient fine-tuning schemes.
@@
Full-Rank $34.06$ (0.360.36G) $25.08$ (0.760.76G) $18.80$ (2.062.06G) $15.56(7.80\mathrm{G})$
GaLore $34.88$ (0.240.24G) $25.36$ (0.520.52G) $18.95$ (1.221.22G) $15.64(4.38\mathrm{G})$
Low-Rank $78.18$ (0.260.26G) $45.51$ (0.540.54G) $37.41$ (1.081.08G) $142.53(3.57\mathrm{G})$
LoRA $34.99$ (0.360.36G) $33.92$ (0.800.80G) $25.58$ (1.761.76G) $19.21(6.17\mathrm{G})$
ReLoRA $37.04$ (0.360.36G) $29.37$ (0.800.80G) $29.08$ (1.761.76G) $18.33(6.17\mathrm{G})$
SUMO 34.26 ($\mathbf{0.23}$G) 24.87 ($\mathbf{0.51}$G) 18.69 (1.161.16G) 14.68 (3.843.84G)
Training Tokens
1.11.1B 2.22.2B 6.46.4B 13.1B
@@
To evaluate the performance of our method on a complex reasoning task, we utilize the GSM8K dataset [47] to test systematic generalization. For these experiments, we used a batch size of 32 and 10 epochs for fine-tuning. We present the performance result in Table 5 training Phi-2 (2.7B) model [48], and in Table 5 training LLaMA (3B) model [49]. The results demonstrate that the proposed method significantly improves generalization to out-of-distribution data. We used NVIDIA H200 GPU.
@@
Zero-shot generalization.
Following this, we extend our evaluation to diverse commonsense and reasoning tasks using zero-shot methods, and then conduct the exact evaluation protocol described in Table 4 of [3] (which suggests the Apollo optimizer). The task details are presented in Appendix D.1. To ensure a fair comparison, we repeat the exact experimental setup used for AdamW, APOLLO, and APOLLO-Mini. The results are summarized in Table 6 below.
Table 6: Zero-shot evaluation of LLaMA-350M models pretrained with sequence length 1024 across reasoning tasks (lower is better for perplexity; higher is better otherwise).
Method Memory Perplexity BoolQ RTE HS WG OBQA ARC-E ARC-C PIQA SciQ MathQA Avg.
AdamW 1.37G 16.30 0.4917 0.4693 0.3688 0.5233 0.332 0.3729 0.2449 0.6534 0.609 0.2064 0.4272
APOLLO 0.34G 15.64 0.5373 0.4698 0.3850 0.4925 0.322 0.3788 0.2483 0.6681 0.624 0.2127 0.4406
APOLLO-Mini 0.15G 16.12 0.5376 0.4562 0.3707 0.5217 0.324 0.3758 0.2312 0.6638 0.619 0.2224 0.4374
SUMO 0.18G 15.49 0.5479 0.4709 0.3937 0.5313 0.321 0.3832 0.2496 0.6709 0.623 0.2246 0.4416
As shown, the SUMO-pretrained LLaMA-350M model achieves lower perplexity on average and consistently outperforms on downstream benchmarks.
@@ 5 Discussion
Empirically, SUMO outperforms prior low-rank methods on both fine-tuning and pre-training tasks, achieving greater memory reduction than memory-efficient benchmarks such as Galore. Our theoretical analysis further confirms its superior convergence properties under practical conditions. These findings position SUMO as a simple yet effective alternative to approximate geometric optimizers. Future work may investigate parallel computations for orthogonalization, integrate quantization techniques, and assess the effectiveness of the method in knowledge editing [50] or domain generalization [51, 52].